Calculate the first fundamental form, the second fundamental form, and the
Weingarten matrix of the following surface patches of the unit sphere:
(a) σ(θ, φ) = (cos θ cos φ, cos θ sin φ,sin θ).
(b) σ(u, v) = (sech u cos v, sech u sin v, tanh u).
1
Expert's answer
2020-11-16T07:50:04-0500

We remind that for the parametric surface "r(u,v)" the first fundamental form is (see http://web.mit.edu/hyperbook/Patrikalakis-Maekawa-Cho/node28.html):

"I=ds^2=Edu^2+2Fdudv+Gdv^2," where "E=(r_u,r_u), F=(r_u,r_v)," "G=(r_v,r_v)" .

The second fundamental form is: "II=Ldu^2+2Mdudv+Ndv^2" , where: "L=-(r_u,\text{N}_u),M=-(r_v,\text{N}_u),N=-(r_v,\text{N}_v)" .

The normal vector is: "\text{N}=r_u\times r_v=\begin{vmatrix}
 i & j& k \\
 (r_u)_1 & (r_u)_2& (r_u)_3\\
(r_v)_1 & (r_v)_2& (r_v)_3
\end{vmatrix}" , where by "()_j" we denoted the respective coordinates of the vectors.

The Weingarten matrix is the matrix of the shape operator of surface s with respect to variables u and v (see https://resources.wolframcloud.com/FunctionRepository/resources/WeingartenMatrix).

It is given by"\begin{pmatrix}
 \frac{MF-LG}{EG-F^2} & \frac{LF-ME}{EG-F^2} \\
 \frac{NF-MG}{EG-F^2} & \frac{MF-NE}{EG-F^2}
\end{pmatrix}"


a). "r_{\theta}=(-sin\,\theta\,cos\varphi,-sin\,\theta\,sin\varphi,cos\,\theta),"

"r_{\varphi}=(-cos\,\theta\,sin\varphi,cos\,\theta\,cos\varphi,0),"

"E=(r_{\theta},r_{\theta})=sin^2\theta\,cos^2\,\varphi+sin^2\theta\,sin^2\,\varphi+cos^2\theta=1" ,

"F=(r_{\theta},r_{\varphi})=sin\,\theta\,\,cos\,\theta\,cos\varphi\, sin\varphi-sin\,\theta\,\,cos\,\theta\,cos\varphi\, sin\varphi=0,"

"G=(r_{\varphi},r_{\varphi})=cos^2\,\theta\,sin^2\varphi+cos^2\,\theta\,cos^2\varphi=cos^2\,\theta."

"I=d\theta+cos^2\theta\,d\varphi" .

"\text{N}=\begin{vmatrix}
 i & j &k \\
 -sin\,\theta\,cos\varphi&-sin\,\theta\,sin\varphi&cos\,\theta\\
-cos\,\theta\,sin\varphi&cos\,\theta\,cos\varphi&0
\end{vmatrix}=i\begin{vmatrix}
 -sin\,\theta\,sin\varphi&cos\,\theta\\
cos\,\theta\,cos\varphi&0
\end{vmatrix}-j\begin{vmatrix}
 
 -sin\,\theta\,cos\varphi&cos\,\theta\\
-cos\,\theta\,sin\varphi&0
\end{vmatrix}+k\begin{vmatrix}
 
 -sin\,\theta\,cos\varphi&-sin\,\theta\,sin\varphi\\
-cos\,\theta\,sin\varphi&cos\,\theta\,cos\varphi
\end{vmatrix}="

"=-i\,cos^2\theta\,cos\,\varphi-j\,cos\,\theta\,sin\,\varphi-ksin\,\theta\,cos\,\theta=(-\,cos^2\theta\,cos\,\varphi,-\,cos\,\theta\,sin\,\varphi,-sin\,\theta\,cos\,\theta),"

The second fundamental form is:

"II=Ldu^2+2Mdudv+Ndv^2,"

where "L=-(r_u,\text{N}_u),M=-(r_v,\text{N}_u),N=-(r_v,\text{N}_v)"

The Weingarten matrix is: "\begin{pmatrix}
 \frac{MF-LG}{EG-F^2} & \frac{LF-ME}{EG-F^2} \\
 \frac{NF-MG}{EG-F^2} & \frac{MF-NE}{EG-F^2}
\end{pmatrix}"

b). "r_{u}=(-\frac{sinh\,u}{cosh^2\,u}cos\,v,-\frac{sinh\,u}{cosh^2\,u}sin\,v,\frac{1}{cosh^2u}),"

"r_{v}=(-sech\,u\,sin\,v,sech\,u\,cos\,v,0),"

"E=(r_u,r_u)=\frac{sinh^2\,u}{cosh^4\,u}cos^2\,v+\frac{1}{cosh^4\,u}=\frac{cos^2v}{cosh^2u}" ,

"F=0, G=(r_v,r_v)=\frac{1}{cosh^2u};"

"I=\frac{cos^2v}{cosh^2u}du+\frac{1}{cosh^2u}dv" .

"\text{N}=\begin{vmatrix}
 i & j &k \\
 -\frac{sinh\,u}{cosh^2\,u}cos\,v&-\frac{sinh\,u}{cosh^2\,u}sin\,v&\frac{1}{cosh^2u}\\
-sech\,u\,sin\,v&sech\,u\,cos\,v&0
\end{vmatrix}=i\begin{vmatrix}
 -\frac{sinh\,u}{cosh^2\,u}sin\,v&\frac{1}{cosh^2u}\\
sech\,u\,cos\,v&0
\end{vmatrix}-j\begin{vmatrix}
 -\frac{sinh\,u}{cosh^2\,u}cos\,v&\frac{1}{cosh^2u}\\
-sech\,u\,sin\,v&0
\end{vmatrix}+k\begin{vmatrix}

 -\frac{sinh\,u}{cosh^2\,u}cos\,v&-\frac{sinh\,u}{cosh^2\,u}sin\,v\\
-sech\,u\,sin\,v&sech\,u\,cos\,v
\end{vmatrix}="

"\text{N}=i\frac{cos\,v}{cosh^3u}+j\frac{sin\,v}{cosh^3u}-k\frac{sinh\,u}{cosh^3\,u}" .

The second fundamental form is:

"II=Ldu^2+2Mdudv+Ndv^2,"

where "L=-(r_u,\text{N}_u),M=-(r_v,\text{N}_u),N=-(r_v,\text{N}_v)."

The Weingarten Matrix is: "\begin{pmatrix}
 \frac{MF-LG}{EG-F^2} & \frac{LF-ME}{EG-F^2} \\
 \frac{NF-MG}{EG-F^2} & \frac{MF-NE}{EG-F^2}
\end{pmatrix}" .